### v2

# NuSTAR discovers a long type-I X-ray burst from the clocked burster GS 1826-24

Abstract
The source GS 1826-24 is a neutron star low mass X-ray binary known as the “clocked burster” because of its extremely regular bursting behavior. We report on the detection of a long type-I X-ray burst from this source. We perform a detailed spectroscopic analysis of the long X-ray burst, lasting for s, seen in the NuSTAR observation carried out on 2022 September. The persistent emission is well described by an absorbed thermal Comptonization model nthcomp, and the source exhibits a soft spectral state during this observation. The observed burst exhibits a rise time of s and a decay time of s. The time-resolved spectroscopy of the burst shows a significant departure from a pure thermal spectrum and is described with a model consisting of a varying-temperature blackbody plus an evolving persistent emission component. We observe a significant enhancement in the persistent emission during the burst. The enhancement of the pre-burst persistent flux is possibly due to Poynting-Robertson drag or coronal reprocessing. At the peak of the burst, the blackbody temperature and the blackbody emitting radius reached a maximum of keV and km, respectively. The peak flux () during the burst is ergs cm-2 s-1, which corresponds to a luminosity of ergs s-1.
Keywords: accretion, X-ray burst - stars: neutron - X-rays: binaries - stars: individual GS 1826-24
1 Introduction
Type I X-ray bursts are thermonuclear explosions on the surface layers of weakly magnetized neutron stars (NSs) accreting mass from a low-mass companion, which is usually a low-mass star () or a white dwarf (Lewin et al. 1993, Strohmayer and Bildsten 2006). The NS accretes mainly hydrogen and/or helium from its Roche lobe filling companion star. The energy source of X-ray bursts is nuclear burning on the surface of NSs, where accreted hydrogen and helium from a companion star are converted into heavy nuclei (Galloway and Keek, 2021). These bursts last from tens to hundreds of seconds, depending on the ignition depth and the composition of the burning material (Bildsten, 2000). Observed burst profiles show a rapid rise followed by an exponential decay, representing the cooling of the NS surface. Short type- I X-ray bursts, lasting up to tens of seconds, are attributed to the triple- process in a helium-rich fuel. The longer bursts of up to a few hundred seconds are expected from hydrogen-rich or mixed H/He fuels (Fujimoto et al. 1981, Bildsten 2000, Cumming 2003). During some highly energetic bursts, the radiation pressure may reach the Eddington limit, so the burning layer expands, lifting off from the surface, thus leading to the expansion of the photosphere of the NS (Cumming 2003, Kuulkers et al. 2003). During the thermonuclear burning, the supply of the seed photons from the surface of the NS is expected to increase and cool the Comptonized corona around the NS (Maccarone and Coppi 2003, Chen et al. 2013, Ji et al. 2013, Sánchez-Fernández et al. 2020). The burst X-ray spectra are usually fit with a blackbody (kT peaking at 2–3 keV; Swank et al. 1977) that cools down during the burst decay. To model the continuum burst emission, one usually assumes that the persistent emission remains unchanged during the burst, and one subtracts it off as part of the background (see, e.g., Kuulkers et al. 2003). However, Worpel et al. (2013) proposed that the persistent emission intensity is variable, thus involving a multiplicative factor to the persistent emission (assuming an unchanged spectral shape) for fitting the spectrum during the burst. In this varying persistent flux model, the value of seems to be proportional to the intensity of bursts.
The source GS 1826-24 is a typical NS Low Mass X-ray Binary (LMXB), located at a distance of 5.7 kpc (Chenevez et al., 2016). The source was initially classified as a blackhole candidate because of its similarity with Cyg X-1. Later, the detection of the three type-I X-ray bursts with BeppoSAX established that the compact object in this system is a weakly magnetized NS (Ubertini et al., 1997). This source GS 1826-24 was discovered as a new transient by Ginga in 1988 September (Tanaka, 1989). It has been X-ray bright and exhibiting regular type-I X-ray bursts since its discovery. The source GS 1826-24 is commonly known as the “clocked burster” because of its extremely regular bursting behavior with a burst recurrence time close to 6 hours (Ubertini et al. 1999, Galloway et al. 2004). GS 1826-24 is a unique X-ray burster that shows remarkable agreement with theoretical models of recurrence times, energetics, and lightcurves. Galloway et al. (2004) studied 24 bursts observed from GS 1826-24 by the RXTE between 1997 November and 2002 July and measured a relationship between persistent X-ray flux and burst recurrence time. The source GS 1826-24 remained in the persistent hard spectral state from 1996 until 2014, and the spectral state was characterized by a dominant power-law component (Galloway et al., 2008, Cocchi et al., 2011, Asai et al., 2015, Rodi et al., 2016). The source exhibited a clocked burster state with burst recurrence times of 3 to 6 hours between 1997 and 2007, as reported by Thompson et al. (2008). Around 2013, the source gradually brightened and transitioned to the soft state after 2014 (Fig. 12 in Asai et al. 2022).
The source was not active in its clocked burster state, and the pattern of regular bursting behaviour has been broken, with the transition to weaker, irregular bursts accompanied by a soft spectral state, in 2014 June (Chenevez et al., 2016). Shortly after this soft episode, the source returned to its hard state again. According to observations by Swift/BAT and MAXI, the intensity of GS 1826-238 gradually decreased and transitioned back to the hard state (Chenevez et al., 2016, Ji et al., 2018). After 2016, this source transfers into intermediate and soft states (Ji et al., 2018). There have been 15 NuSTAR StrayCat observations of GS 1826-24 between 2014 February and 2021 November, and a transition of the source from the island atoll state (hard) to a banana branch (soft) has been observed during this time interval (Yun et al., 2023). Again, approximately after one year, NuSTAR observed the source in 2022 September in an extended soft spectral state, and here we present the analysis of this NuSTAR observation. Presumably, the change in the observed behavior is correlated to a change in the mass accretion rate onto the NS (Hasinger and van der Klis, 1989). Recent observation suggests that GS 1826-238 may have entered again a clocked burster state for the first time in 10 years with a shortest burst recurrence period of approximately 1.6 hours (Iwata et al., 2025).
Chenevez et al. (2016) detected the Eddington limited type-I X-ray burst and derived the distance of the source to be kpc, assuming an isotropic emission. An excess in the soft flux and a shortage in the hard X-ray flux was reported during the type-I X-ray burst observed in this source (Ji et al. 2014, Sánchez-Fernández et al. 2020). The source is also known for the long burst tails, lasting s, which were powered by rp-process hydrogen burning (Galloway et al., 2004). The inferred accretion rate varies little on short timescales, and is within the range (Galloway et al. 2008, Chenevez et al. 2016). The NICER observations of this source revealed 7–9 mHz oscillations with fractional rms 2 per cent at 6 keV (Strohmayer et al., 2018). The first superburst (lasted for at least 3 hours) from GS 1826–24 was detected by the MAXI/GSC on 2018 February (Iwakiri et al., 2018).
In the present work, we have carried out a detailed time-resolved spectral analysis of the long type-I X-ray burst observed during the NuSTAR observation. We have also performed the spectral analysis of the persistent emission. An enhancement of the persistent emission has been observed during the burst. We discuss this in detail. This paper is structured as follows. In Section 2, we describe the observations and data reduction used in this study, and in Section 3, we discuss the light curve of the burst. Section 4 describes the spectral analysis techniques adopted for the persistent and burst emission. Finally, we discuss the observational findings and their implications in Section 5.
2 observation and data reduction
The Nuclear Spectroscopic Telescope Array (NuSTAR) is a hard X-ray (3–79 keV) focusing telescope with two identical, co-aligned telescopes, FPMA and FPMB. Its effective area peaks at cm2 (adding up the two modules) around 10 keV, and its energy resolution at 10 keV is 400 eV (Harrison et al., 2013). NuSTAR observed the source on 12th September 2022 (obs ID: 30801022002) for a total of ks. The current analysis is based only on this observation.
We processed the NuSTAR data using the data analysis Software NuSTARDAS v2.1.5, which is distributed with HEASOFT v6.35.1, and used the latest calibration files v20250415 available during the analysis. We used the task nupipeline v0.4.12 to generate the calibrated and screened event files. A circular extraction region of radius was used to study the persistent and burst spectrum. To collect background events, we selected a region of similar size from the same chip but in a position away from the source. Spectra and light curves were extracted using the FTOOL nuproducts from the FPMA and FPMB. During this observation, we created good time intervals (GTIs) for the persistent emission and the observed type-I X-ray burst. For the time-resolved spectroscopy, we created seven segments along the burst profile and built the corresponding GTI files. We then used them to extract the burst spectra for each segment. We found that the background dominates the signal at energies over keV.
3 Light Curve
In the left panel of Figure 1, we show the NuSTAR/FPMA light curve of the source in the keV energy band. The average count rate during the NuSTAR observation was counts s-1. One type-I thermonuclear X-ray burst is present at about th ks (MJD ) during the NuSTAR observation (see left and middle panel of Figure 1). From this particular NuSTAR orbit, a change in the persistent flux has been observed (see right panel of Figure 1). The X-ray burst has a typical sharp rise and near-exponential decay. The observed burst has a long tail, lasting s, and the peak of the burst seems to reach approximately counts s-1, about a factor of above the level of the persistent emission. Considering the rising and exponential decay time, the type-I burst lasts about s. We extracted the light curve (see Figure 2) of the burst in different energy bands. We performed the time-resolved spectral analysis of the burst to investigate the spectral properties in detail. For the persistent emission spectrum, we took a time interval of around s just before the rise of the burst and a time interval of ks before the burst, but from a different NuSTAR orbit.
4 spectral analysis
The persistent and the burst spectra are analyzed with the X-ray spectral package XSPEC (Arnaud, 1996). For the persistent emission, we model the NuSTAR FPMA and FPMB spectra simultaneously between to keV energy band. To account for cross-calibration of the two instruments, FPMA and FPMB, a constant multiplication factor constant was included in the modeling. The value of constant for FPMA was fixed to , allowing it to vary for the FPMB. We used the TBabs model to account for interstellar absorption along the line of sight with the wilm abundances (Wilms et al., 2000) and the vern (Verner et al., 1996) photoelectric cross-section. Spectral uncertainties are quoted at the confidence level.
4.1 Persistent spectral analysis
We investigate the persistent emission before the burst to quantify the spectral shape and to measure the persistent flux for comparison with the burst spectrum. We generate a pre-burst spectrum of the source from two different time intervals: (1) a time interval of 2 ks (from the elapsed time 41st ks to 43rd ks) has been chosen before the burst but from a different NuSTAR orbit (2) another time interval of s has been chosen just before the burst onset depending on the data available in the same NuSTAR orbit. After generating the keV NuSTAR FPMA and FPMB persistent spectrum, we tried to fit them simultaneously with an absorbed cutoff power law model TBabs*cutoffpl. Due to the lack of coverage below 3 keV, we fixed the hydrogen column density at cm-2 (Dickey and Lockman, 1990). We found that both persistent emission spectra are well described by the model tbabs*cutoffpl with and for the two persistent spectra, respectively. The persistent emission is characterized by a power law component with a photon index of and the cutoff energy . This power law component can be explained as the result of inverse Compton scattering in the corona. However, the direct disc emission is not spectrally detected due to the low energy limitation of NuSTAR.
To describe the persistent emission more precisely, we employed a thermal Comptonization model nthcomp (Zdziarski et al., 1996), which attempts to simulate the upscattering of photons through the corona from a seed spectrum parameterized either by the inner temperature of the accretion disc or the NS surface/boundary layer. We then tested the physically-motivated Comptonization model nthcomp, setting the photon seed input to a disc blackbody. This model TBabs*nthcomp also describes the continuum emission very well with and for two persistent spectra, respectively (see Figure 3). We estimated the electron temperature at with photon index . The seed photon temperature is found to be low , indicating the difficulties in detecting the soft disc blackbody component arising from the accretion disc. The electron temperature of the corona, , is roughly comparable to the observed cutoff energy ( keV) of the spectrum. The best-fit parameters for the persistent spectrum are quoted in Table 1.
| Component | Parameter (unit) | Persistent (2ks) | Persistent (400s) |
| Constant | FPMB (wrt FPMA) | ||
| tbabs | () | (f) | (f) |
| cutoffpl | |||
| Norm | |||
| nthcomp | |||
| Norm | |||
| cflux | |||
Flux values of nthcomp component have been calculated in the keV energy band in the unit of ergs cm-2 s-1. The parameter inp-type of nthcomp was fixed to (disk photon).
4.2 Burst spectral analysis
We performed time-resolved spectroscopy of the long burst to investigate the evolution of the spectral parameters throughout the burst interval. We divided the entire burst interval into seven segments of different time intervals after the burst onset to cover the rising phase, peak, and long tail of the burst. The time intervals were chosen to be between s (segment 1, S1), s (S2), s (S3), s (S4), s (S5), s (S6), and s (S7) from the onset of the burst light curve (see Figure 2), following Sánchez-Fernández et al. (2020). After s, the light curve of the burst gradually flattens and eventually reaches the persistent level around s. We extracted the spectrum of each segment separately with a minimum of 100 counts in each spectral bin in the FPMA spectrum. To fit the burst spectra, we used an absorbed blackbody model TBabs*bbodyrad with fixed absorption cm-2. We found that a single blackbody radiation could not fit the burst spectrum. Although the burst spectra are dominated by a blackbody, excesses are evident at both low (soft X-ray excess) and high energy (hard X-ray deficit), see the upper panel of Figure 4. All the spectra extracted from different segments showed similar characteristics.
To take into account the residuals of the blackbody burst fit, we employed the enhanced persistent emission model. Initially, we performed the fit assuming that the persistent emission spectrum does not change during the burst, corresponding to the standard burst fitting technique (e.g. Kuulkers et al. 2002, Galloway et al. 2008). This model failed to fit the observed hard X-ray deficit and the soft X-ray excess for all the spectra. Later, we assumed that the amplitude of the persistent emission can change during the burst (Worpel et al., 2013) and adopted the following best-fit model composed of TBabs*(bbodyrad+ nthcomp). The blackbody and nthcomp models account for the X-ray burst and persistent emissions during the burst, respectively. The parameters of the nthcomp were fixed to the best-fit values of the preburst persistent emission taken from s time interval (reported in Table 1). The parameter is a free scaling factor that can be used to account for the variation of the persistent emission. If , this means that the amplitude of persistent emission during the burst is equal to the value before the burst. This model improved the fit and provided a reasonable description of all the burst spectra (see bottom panels of Figure 4). We obtained an acceptable fit for all the intervals of the burst with the model TBabs*(bbodyrad+nthcomp), and it constrained all the free parameters simultaneously. We found that the best-fit values of is greater than for all segments, suggesting that the persistent emission typically increases during the burst. The best-fit parameter values for individual segments are quoted in Table 2. We used the XSPEC model cflux to determine the unabsorbed fluxes of different spectral components of the time-resolved spectra in the energy range keV during the burst (see Table 2).
| Component | Parameter (unit) | S1 (20s) | S2 (20s) | S3 (30s) | S4 (40s) | S5 (50s) | S6 (100s) | S7 (200s) |
| tbabs | (f) | (f) | (f) | (f) | (f) | (f) | (f) | |
| Mult. factor | ||||||||
| cflux | ||||||||
| bbodyrad | (keV) | |||||||
| Norm | ||||||||
| cflux | ||||||||
| cflux | ||||||||
In unit of . Flux values of different spectral component during the burst have been calculated in the keV energy band in the unit of ergs cm-2 s-1. The parameters of nthcomp were fixed to the persistent values (see Table 1).
We tested other scenarios to fit the burst spectra apart from the model. First, we attempted to fit the burst spectra by the model TBabs*thcomp*bbodyrad, considering that the burst photons could be affected by the corona/boundary layer (Chen et al., 2022). Here, thcomp (Zdziarski et al., 2020) is a more accurate version of nthcomp, and the parameter values of this model are extracted from the persistent fit. However, this model failed to fit the residuals observed at low and high energies and provided a fit to the burst spectra that was significantly worse compared to the previous model. Second, we tried the model TBabs*(bbodyrad+powerlaw+nthcomp) to fit the burst spectrum, in which the nthcomp parameter values were fixed to the persistent emission (Jaisawal et al., 2025). However, the current data failed to constrain all the free parameters simultaneously and could not produce conclusive results. Third, to fit the burst spectra, another combination of models like TBabs*(bbodyrad+
relxillNS+nthcomp) can be tested in which relxillNS accounts for the reflection of the burst photons from the accretion disk (Lu et al., 2023). However, we could not implement this model because of the short exposure time of each segment of the burst spectrum and the lack of a priori knowledge of some important disk-related parameters. The above-mentioned facts restricted us from discussing these models.
5 Discussion
We performed a spectroscopic analysis of the long X-ray burst, observable for s, seen in the NuSTAR observation of the source GS 1826-24 to investigate the properties of the burst in detail. We first considered the spectral studies of the persistent emission of the source. We extracted persistent spectrum in the keV range from a time interval of 400 s just before the onset of the burst and a time interval of 2 ks before the burst, but from a different NuSTAR orbit. We modeled the persistent spectra using an absorbed cutoff power law (cutoffpl) or a thermal Comptonization model (nthcomp). Both models yielded a reasonably good description of the persistent emission in both cases. During this observation keV persistent flux was ergs cm-2 s-1, which correspond to a luminosity of ergs s-1 (assuming a source distance of kpc). This luminosity is roughly of the Eddington luminosity (where ergs s-1 for a NS). The inferred persistent flux is consistent with the intermediate and soft state bolometric flux of the source (Ji et al., 2018, Sánchez-Fernández et al., 2020, Grefenstette et al., 2025). The local accretion rate per unit area can be calculated from the persistent luminosity following Galloway et al. (2008), which is g cm-2 s-1 for a NS of mass and radius 10 km in our case. Thus, the accretion rate onto the source was modest and exhibited a soft spectral state during this observation.
We extracted seven separate spectra along the burst profile, in the intervals 20 s, 20 s, 30 s, 40 s, 50 s, 100 s, and 200 s from the burst onset. We initially modeled the burst spectra using a simple absorbed blackbody model, but excesses are visible at low and high energies. To fit the time-resolved spectra of the burst properly, we used a model consisting of a blackbody plus a variable persistent emission component. We accounted for the excess flux in X-ray burst emission using the model of Worpel et al. (2013). Along the burst profile, the blackbody temperature decreases from the peak value keV to keV during the tail of the burst. The blackbody emitting radius () is highest km at the peak of the burst and becomes smaller as the flux decreases, but not in a regular fashion (see Figure 5). We did not observe any indication of photospheric radius expansion (PRE) at the peak of the burst from the time-resolved spectroscopy. The peak flux () during the burst is ergs cm-2 s-1. This flux corresponds to a luminosity of ergs s-1, which is roughly of the , indicating that the flux at the peak of the burst did not reach the Eddington limit. Additionally, we found that the best-fit values of are greater than , suggesting that the persistent emission typically increases during the burst. The values rise at the burst start and quickly reach the maximum () within seconds, then they decrease slowly and return to around unity during the cooling tail (see top panel of Figure 5). This indicates that the persistent emission is rapidly enhanced due to the X-ray burst radiation and then returns to the preburst level during the cooling tail. To verify this, we further extracted a spectrum from a segment (after S7) of time interval s ( s) and fit in the same way as earlier for the spectra of segments S1…S7. The fit returned value of , thus confirming the above-mentioned fact. The enhancement in is perfectly consistent with the results obtained by Ji et al. (2015) after analyzing burst spectra seen by RXTE. We determined the unabsorbed keV blackbody flux (), the persistent flux (), and showed their variation during the burst in the Figure 6. Around the burst peak, the blackbody flux supersedes the persistent flux; later on, the persistent flux starts dominating.
Our spectral analysis of the burst does not agree with the traditional assumption that the persistent, Comptonized emission does not vary during an X-ray burst. The time-resolved burst spectra from GS 1826-24 showed a significant departure from a pure thermal spectrum. During the entire burst interval, we observed an excess of soft and hard emission over the simple blackbody model. At the peak of the burst, the hard excess is less pronounced compared to the decaying phase (see top panel of Figure 4). We used a variable persistent emission to account for the excess flux in the X-ray burst emission. For this source, a drop in the hard X-ray emission, together with a soft X-ray excess with respect to the burst blackbody emission, has been observed earlier by Sánchez-Fernández et al. (2020) from the INTEGRAL and XMM-Newton observations. From the RXTE observations, hard X-ray shortages and enhancements of the persistent emission at soft X-rays during the burst have been observed from this source (Ji et al., 2015). However, The detection of excess burst emission has been observed for several other NS LMXBs: 4U 1820-30 (Keek et al., 2018b), Aql X-1 (Keek et al., 2018a), SAX J1808.4-3658 (in’t Zand et al. 2013, Bult et al. 2019), 4U 1608-52 (Jaisawal et al. 2019, Chen et al. 2019), 4U 1636-536 (Güver et al. 2022, Kashyap et al. 2022), 4U 1728-34 (Bhattacharyya et al., 2018), due to the increased sensitivity afforded by NICER, AstroSat, and Insight-HXMT. This flux excess may indicate interactions between the burst emission and the accretion flow surrounding the NS. There may be various modes of interaction, but three processes are commonly considered to explain the excess flux. First, the burst emission may be reprocessed in the surface layers of the accretion disk and reflected into the line of sight (Ballantyne and Strohmayer, 2004). In this case, the magnitude of the reflection component should approximately follow the burst intensity. Second, these excesses could be due to enhanced accretion rate onto the stellar surface through Poynting–Robertson (PR) drag (Walker, 1992), which corresponds to the change in persistent emission normalization (Worpel et al., 2013). Third, The stellar surface could be covered by a Comptinizing medium, such that our view of the NS is partially obscured (Kajava et al., 2014), thereby the flux excesses may be a result of the Comptonization (Keek et al., 2018a) or scattering processes in the atmosphere (Romani, 1987).
It is important to understand how much of the observed X-ray emission is directly due to the thermonuclear process on the stellar surface and how much is added by the interaction with the stellar atmosphere. We, therefore, calculated the total flux along the X-ray burst profile and the contribution of the blackbody and the variable persistent emission separately. We found that the blackbody flux increases during the peak of the burst, and the contribution of the persistent emission is minimal (accounts for some of the total flux) during this interval.
It is expected because when an X-ary burst occurs in the NS envelope, the number of soft seed photons increases dramatically, enhancing the blackbody flux. However, the enhancement of the persistent flux is observed outside of the peak of the burst, and its contribution increases from to of the total flux during this decay phase (see Figure 6). We, therefore, observed a significant enhancement in the persistent emission during the burst. However, the underlying mechanism of such excess is not well understood. It has been observed that if the X-ray burst irradiates the accretion disk, it simultaneously results in both reflection and PR drag (Fragile et al., 2020). We did not observe any reflection signature in the burst spectrum, which indicates that the first possibility of the excess emission during the burst may be ruled out. Thus, the enhancement in the persistent emission may be because of the PR effect, which temporarily increased the accretion rate onto the stellar surface (Walker, 1992) or some other reprocessing of burst photons by the corona or the boundary layer. However, the data for the X-ray burst presented here are not sensitive enough to distinguish between these possibilities clearly. On the contrary, an apparent shortage of the persistent emission during the burst at hard X-rays (above 30 keV) indicates cooling of a corona (Sánchez-Fernández et al., 2020).
The total duration of the burst, i.e., the time to evolve away from and return to the persistent state, was s in the NuSTAR FPMA/FPMB lightcurves. The burst has characteristics of both short-duration soft-state bursts and long-duration hard-state bursts (Yun et al., 2023, Grefenstette et al., 2025). During this burst duration, the intensity of the source increased by at least (at the beginning and end point of the burst) compared to its persistent intensity level. We found that at about seconds after the burst onset, the flux reached of its peak flux. The burst exhibited a moderate rise time of s, which is the time spent between the start of the burst and the point at which of the peak burst intensity is reached. The bolometric flux ( keV) during the burst is ergs cm-2 s-1 which has been used to calculate the burst fluence (). The value of was found to be ergs cm-2. The corresponding decay time of the burst is s. The measured fluence of the burst corresponds to a net burst energy release of erg (assuming a distance of kpc). The ignition depth () is related to the energy release during the burst by the equation , where is the energy release during nuclear reaction (Cumming, 2003). To calculate , we take an energy release of burning He to Ni of erg g-1 (including losses due to neutrino emission), and assume a NS mass of and radius of km, which give a gravitational redshift of . We found an ignitation depth g cm-2. At an accretion rate of g cm-2 s-1, the recurrence time corresponding to a column depth of g cm-2 is days. However, this can not be compared to the earlier recurrence times as the current burst is observed when the source was not in an active clocked burster state.
To understand the burst fuel composition, one can estimate the factor, which is the ratio of the persistent fluence between bursts to the total burst fluence of the burst, using the formula . In our case, there is no scope to calculate from the observation, as we see only a single burst in this observation. Therefore, we verified the burst fuel composition in another way by determining the local accretion rate (Galloway et al., 2008). A convenient unit of accretion rate is the Eddington rate. We, therefore, expressed the local accretion rate () as a fraction of the local Eddington accretion rate (). We assumed the local accretion rate g cm-2 s-1, then we obtained as a fraction of for (for solar abundance) and (for pure helium). We found that and for and , respectively. The value of less than for supports the idea that these bursts were ignited in a helium-rich environment (see e.g., Galloway and Keek 2021) and such accretion rate is expected based on the burst behavior (Galloway et al., 2008). Despite the uncertainty in the fuel composition, the properties of the bursts (i.e., longer burst duration) from GS 1826-24 strongly suggest burning of mixed H/He fuel by unstable He ignition (Fujimoto et al., 1981). However, to study various burst properties in great detail, we need an uninterrupted observation of this source with broadband X-ray coverage.
6 Data availability
This research has made use of data obtained from the HEASARC, provided by NASA’s Goddard Space Flight Center. The NuSTAR observational data set used in this work (Obs. IDs ), dated September 12, 2022, is in the public domain and is available on NASA’s website https://heasarc.gsfc.nasa.gov.
7 Acknowledgements
We want to thank the anonymous referee for his/her feedback and suggestions, which have significantly enhanced the quality of this paper. We would also like to thank Dr. Biplab Raychaudhuri for the productive discussions. This research has made use of data and/or software provided by the High Energy Astrophysics Science Archive Research Centre (HEASARC). This research also has made use of the NuSTAR data analysis software (NuSTARDAS) jointly developed by the ASI Space Science Data Center (SSDC, Italy) and the California Institute of Technology (Caltech, USA). ASM would like to thank Inter-University Centre for Astronomy and Astrophysics (IUCAA) for their facilities extended to him under their Visiting Associate Programme.